# Research Paper

### v1

@@ 3.3 External data sets
For low redshift lags associated with the H line, we also supplement our data with the external lags utilised in the analysis of OzDES-Malik_2023. These include lags from: the Lick AGN Monitoring Project (LAMP) (HBETA_U_2022), a low redshift () study of statistically diverse AGN; the SEAMBH survey, which targets high luminosity sources known to produce lower reverberation lags (HBETA_Du_2016; HBETA_Du_2018; HBETA_Hu_2021); and a collection of low redshift sources collated by HBETA_Bentz_2013, coupled with similar sources from HBETA_Fausnaugh_2017; HBETA_Bentz_2009; HBETA_Bentz_2013; HBETA_Bentz_2014; HBETA_Bentz_2016a; HBETA_Bentz_2016b and HBETA_Bentz_2023. Also included are a number of additional sources from HBETA_Zhang_2019; HBETA_Li_2021; HBETA_Lu_2016; HBETA_Pei_2014; HBETA_Rakshit_2019. Sources prior to 2019 and not associated with the SDSS collaboration are drawn from a convenient collation by HBETA_Martinez-Aldama_2019.
@@
In this work, as with past OzDES RM papers, we use the dispersion-based definition of line width as our measure of velocity rather than the FWHM, and default to this measurement for literature measurements where available in the source paper. In making the decision between mean-spec and RMS-spec, we defer to the preferred method of the source paper, e.g. mean spectrum for sources from HBETA_Bentz_2013, RMS spectra for sources from SDSS-Shen_2023 etc. For source-papers in which only the FWHM line widths are provided, we follow the lead of CIV_Kaspi_2021 and use based on an assumption of a Gaussian line profile.
@@ 4.2.1 Quality Cuts for H Lags
The first cut selects only well constrained lags, the second leverages the robustness of the ICCF to account for spuriously constrained JAVELIN false positives, and the last two cuts ensure that we keep only lags that do not favour the null hypothesis of there being no relation between the two light curves.
@@ 4.2.2 Quality Cuts for MgII Lags
In the SDSS approach, the marginalized lag posterior distribution , as recovered by PyROA, JAVELIN or PyCCF, is attenuated by a weighting function that down-weights lags that correspond to the poorly constrained seasonal gaps. This weighting function is the convolution of two components. The first term, based on the fraction of observations that overlap between the two light curves after shifting by some lag, down-weights ‘off-season’ lags in seasonal gaps where this overlap is small. The second term is the auto-correlation function of the continuum signal (ACFCont), which smooths and widens the ‘on-season’ peaks of the weighting function to account for the continuous nature of the light curve. The entire posterior is then convolved with Gaussian smoothing kernel of width days (equation 5):
@@
OzDES-Yu_2023 also includes quality cuts modelled after the more rigorous simulation-work of OzDES-Penton_2021, in which simulations of mock DRW signals are used to characterize false positive rates, and quality cuts tuned to remove suspicious sources. These cuts, designed specifically for unweighted JAVELIN posterior distributions and OzDES-like observations, require:
@@ 4.2.3 Quality Cuts for CIV Lags
OzDES tracked high redshift AGN with CIV lines visible for reverberation mapping. These distant sources represent the earliest and most luminous AGN in the OzDES RM sample, and, owing to the anti-correlation of AGN luminosity and optical variability (MacLeod_2010), the sources with the weakest AGN variability. These sources are also observed with the shortest rest-frame time window due to higher time dilation at these higher redshifts. In combination, lag recovery for the CIV sample is a difficult process with an outsized risk of false positives if approached naively. For CIV RM of the OzDES sample, OzDES-Penton_2021 provide a series of quality cuts to maximize the reliability of constraints and minimize the rate of false positives, tuned on mock simulations of OzDES-like data. Under these cuts, sources are sorted into quality levels of ‘bronze’, ‘silver’, ‘gold’, or are rejected completely based on the properties of their marginalized lag posterior distribution and the lag recovery results from the ICCF method. These cuts are:
@@ 4.2.4 SDSS Sub-Sampling Criteria
By contrast to SDSS, OzDES restrict our lag search range to only the physically reasonable domain and apply stringent cuts to minimise the effect of aliasing on each individual source. This produces a more restricted sample with fundamentally different statistical properties, which can be seen in the tension between the SDSS and OzDES constraints. In particular, restricting the SDSS sample to OzDES-like quality cuts significantly lowers scatter about the MgII relationship (discussed in detail in Section 6.1.2). To make joint fits to the relation using both the SDSS and OzDES data sets, we required a set of lags with relatively homogeneous selection properties. To that end, we derive an alternate set of lags from the published SDSS data of SDSS-Shen_2023 that differ from their sample in two ways: firstly, in that we use the lags as recovered from JAVELIN instead of their choice of PyROA, and secondly that we apply a number of quality cuts to this alternate sample to bring it more in line with the rationale of the OzDES selections.
@@
For step 2, in cases where the JAVELIN lag peak is strongly constrained, we replace the JAVELIN confidence interval in this test with days for H or days for CIV and MgII, following a similar reasoning to the cuts of OzDES-Penton_2021. After applying these cuts, we retain a sub-sample with H lags, MgII lags and CIV lags from the SDSS sample.
This SDSS-JAVELIN subsample is broadly similar to the published SDSS sample for the H and CIV lags, but prefers a markedly for lower scatter in the MgII relationship. For the MgII lags, these selection criteria drastically reduce the scatter of the lags about their best-fit relationship, and marginally decrease its preferred slope (see Fig. 11(a) for a comparison).
@@ 5.1 Lag measurements
In total, the OzDES reverberation mapping program has produced high quality AGN lags, consisting of H lags (OzDES-Malik_2023), MgII lags (OzDES-Yu_2021; OzDES-Yu_2023) and CIV lags (OzDES-Penton_2025). For each of these sources, we calculate the mass using equation 1 and accretion rate using equation 4, where the monochromatic luminosity uses bolometric corrections of Runnoe_2012 (Table 2 in their paper) and we use the virial factor of from Grier_2013b. Rest-frame lags range from 333 being days, here and elsewhere in the paper. to , sitting within the range of lags recovered by other works. We find that the distribution of lags over redshift is reasonably consistent with existing reverberation mapping results from the literature (Fig. 4, upper left). We note that the distribution of lags changes abruptly as one moves between lines, most notably between the MgII and CIV lines, indicating different lags, and thus radii, for different ions. This is quantified in Section 7. These RM results, as well as the redshifts and luminosities for the relevant sources, are listed in Table LABEL:tab:Results_All_Hbeta for our H sources, Table LABEL:tab:Results_All_MgII for MgII sources, and Table LABEL:tab:Results_All_CIV for CIV sources.
@@ 5.2 Black hole masses from direct RM measurements
We broadly see an increase in mass with redshift (Fig. 4 lower left), which is expected, as the AGN at higher redshifts are selected to be more luminous (see, for example, Fig. 1 in OzDES-Malik_2024), and this selection effect will dominate any evolutionary effects. We similarly see the expected general increase of more luminous AGN being more massive, but are unable to infer a clear power-law relationship (Fig. 4 lower right).
It should be noted also that the estimated accretion rates for high redshift CIV sources can be extremely high (Fig. 4 upper right), with most direct RM samples sitting above the threshold used by Du_2015 as the division between low and high accretion sources. In Section 6.1.3 we note an apparent flattening of the CIV curve at high luminosity, and, given that SEAMBH associate high accretion rates with low lags (HBETA_U_2022), this offers a possible physical explanation. Within the OzDES sample, there is a weak trend for more luminous, higher-redshift sources to accrete more strongly (black data points, Fig. 4 upper right).
@@ 6 Comparison of New Results With Past Surveys & Tuning of Relationship
Here we examine the relationship using the full array of lags available at the close of this latest generation RM surveys. In addition to lags from OzDES and the final SDSS release of SDSS-Shen_2023, we use a wide array of past H, MgII and CIV lags as outlined in Section 3.3. In Section 6.1, we use a Bayesian approach to constrain the parameters of the relationship for all three reverberating emission lines, and to identify where the results of different surveys are in tension.
@@ 6.1 Comparison of Parameter Constraints
In this section we fit the relationship using OzDES data alone and using OzDES data in combination with data from other sources. We note that some surveys are in statistical tension, which indicate unexplored selection effects, and so Table 2 lists the parameters for our chosen ‘primary’ data sets, while the following sub-sections elaborate on alternate relationships with different combinations of literature data and quality cuts.
@@
Here, represents the choice of luminosity units, typically or erg/s, and is chosen to minimize the covariance between offset and slope in the posterior distribution. varies with the choice of time units for the lag, and we here follow the convention of all lags being in days.
@@
All fits are performed using the Bayesian analysis tool NumPyro (numpyro), specifically its No U-Turn (NUTS) sampler, a Hamiltonian Monte Carlo (HMC) method that is well suited to unimodal distributions. We note that this fitting is relatively insensitive to the choice of sampler, as it is unimodal and produces approximately Gaussian contours (e.g. Fig. 11(a)).
| H | |||||||||||||||||
| MgII | |||||||||||||||||
| CIV | |||||||||||||||||
@@ 6.1.1 H Relationship
Fig. 11(a) (a) shows relationship parameters for data from OzDES, SDSS, the Bentz Collection, the SEAMBH Survey, the LAMP survey, and a number of other miscellaneous sources. Shown also are the contours for SDSS lags as recovered from JAVELIN instead of PyROA, and selected according to the criteria outlined in Section 4.2.4. The constraining power is highest for surveys with a high source-count, like SDSS & SEAMBH, or a wide range of luminosities as in the Bentz Collection.
@@
Fig. 11(a) (b) shows the constraints on relationship parameters when combining all datasets that are not in tension with one another. Four such groups are possible:
@@
Where we use a ‘primary’ H grouping elsewhere in the paper, it refers to the first of these groupings.
@@ 6.1.2 MgII Relationship
This yields two mutually consistent groupings, making use of the full SDSS data set and excluding OzDES or including OzDES but restricting the SDSS data. Owing to the complementary correlations of the data sets, both choices have near identical results for the main properties of the relation, but with a slightly higher scatter for the SDSS-based results (Fig. 11(a) (c)). We proceed with the OzDES +SDSS-JAVELIN + Zajacek fit as our primary MgII relation, as it has both the lowest scatter and is derived from the most diverse set of sources.
@@ 6.1.3 Relationship
The tension between the slope of the high and low redshift is manifest in an apparent ‘levelling out’ of the relation at high redshift (Fig. 5, bottom panel). This is easiest to see when plotting the residuals against the ‘high-’ fit (our ‘grouping 1)’). As shown in Fig. 11, high luminosity / redshift sources give unmistakably and consistently higher lags than the lower luminosity / redshift data.
This nonlinear relationship for CIV may have a physical basis, such as distant AGN having higher accretion rates, which have been associated with lower lags (e.g. Du_2015; HBETA_Hu_2021). A simpler explanation lies in describing this as a selection effect: more luminous AGN have longer rest-frame lags, stretched further at high redshift by time dilation, and are therefore more likely to have lags that exceed the length of the survey. This induces a selection effect in which high redshift / luminosity sources are biased downwards from the true population distribution. CIV sources are often the most distant, and consequently the most luminous, and so their observations are the most likely to be hampered by this effect.
@@
| Survey |
| OzDES | ||||||||||||
| SDSS-JAVELIN | ||||||||||||
@@
This model significantly relaxes the constraints from each survey, introducing the freedom for an un-observed fraction of sources to exist above the lag cutoff. This relaxes the tension between the surveys, creating a single parameter-space region of overlap between all four data sources (Fig. 12, top panel). The high- Kaspi lags are dominated by this observational window effect, making this fit match the constraints of the more generic fit achieved when excluding the high- literature lags all together (Fig. 12, bottom panel). It is this fit, with all CIV data accounted for, that we use as our primary fit for single-epoch mass estimates in Section 8.

### v2

@@ 3.3 External data sets
For low redshift lags associated with the H line, we also supplement our data with the external lags utilised in the analysis of OzDES-Malik_2023. These include lags from: the Lick AGN Monitoring Project (LAMP) (HBETA_U_2022), a low redshift () study of statistically diverse AGN; the SEAMBH survey, which targets highly accreting high luminosity sources known to produce lower reverberation lags (HBETA_Du_2016; HBETA_Du_2018; HBETA_Hu_2021); and a collection of low redshift sources collated by HBETA_Bentz_2013, coupled with similar sources from HBETA_Fausnaugh_2017; HBETA_Bentz_2009; HBETA_Bentz_2013; HBETA_Bentz_2014; HBETA_Bentz_2016a; HBETA_Bentz_2016b and HBETA_Bentz_2023. Also included are a number of additional sources from HBETA_Zhang_2019; HBETA_Li_2021; HBETA_Lu_2016; HBETA_Pei_2014; HBETA_Rakshit_2019. Sources prior to 2019 and not associated with the SDSS collaboration are drawn from a convenient collation by HBETA_Martinez-Aldama_2019.
@@
In this work, as with past OzDES RM papers, we use the dispersion-based definition of line width, as measured on the mean spectrum, as our measure of velocity rather than the FWHM, and default to this measurement for literature measurements where available in the source paper. In making the decision between mean-spec and RMS-spec, we defer to the preferred method of the source paper, e.g. mean spectrum for sources from HBETA_Bentz_2013, RMS spectra for sources from SDSS-Shen_2023 etc. For source-papers in which only the FWHM line widths are provided, we follow the lead of CIV_Kaspi_2021 and use based on an assumption of a Gaussian line profile.
@@ 4.2.1 Quality Cuts for H Lags
The first cut selects only well constrained lags, the second leverages the robustness of the Interpolated Cross-Correlation Function (ICCF) to account for spuriously constrained JAVELIN false positives, and the last two cuts ensure that we keep only lags that do not favour the null hypothesis of there being no relation between the two light curves.
@@ 4.2.2 Quality Cuts for MgII Lags
In the SDSS approach, the marginalised lag posterior distribution , as recovered by PyROA, JAVELIN or PyCCF, is attenuated by a weighting function that down-weights lags that correspond to the poorly constrained seasonal gaps. This weighting function is the convolution of two components. The first term, based on the fraction of observations that overlap between the two light curves after shifting by some lag, down-weights ‘off-season’ lags in seasonal gaps where this overlap is small. The second term is the auto-correlation function of the continuum signal (ACFCont), which smooths and widens the ‘on-season’ peaks of the weighting function to account for the continuous nature of the light curve. The entire posterior is then convolved with Gaussian smoothing kernel of width days (equation 5):
@@
OzDES-Yu_2023 also includes quality cuts modelled after the more rigorous simulation-work of OzDES-Penton_2021, in which simulations of mock DRW signals are used to characterise false positive rates, and quality cuts tuned to remove suspicious sources. These cuts, designed specifically for unweighted JAVELIN posterior distributions and OzDES-like observations, require:
@@ 4.2.3 Quality Cuts for CIV Lags
OzDES tracked high redshift AGN with CIV lines visible for reverberation mapping. These distant sources represent the earliest and most luminous AGN in the OzDES RM sample, and, owing to the anti-correlation of AGN luminosity and optical variability (MacLeod_2010), the sources with the weakest AGN variability. These sources are also observed with the shortest rest-frame time window due to higher time dilation at these higher redshifts. In combination, lag recovery for the CIV sample is a difficult process with an outsized risk of false positives if approached naively. For CIV RM of the OzDES sample, OzDES-Penton_2021 provide a series of quality cuts to maximise the reliability of constraints and minimise the rate of false positives, tuned on mock simulations of OzDES-like data. Under these cuts, sources are sorted into quality levels of ‘bronze’, ‘silver’, ‘gold’, or are rejected completely based on the properties of their marginalised lag posterior distribution and the lag recovery results from the ICCF method (examples of gold and bronze CIV sources can be seen in Figure LABEL:fig:_CIV_example_curves). These cuts are:
@@ 4.2.4 SDSS Sub-Sampling Criteria
By contrast to SDSS, OzDES treats this lag search range as a true prior, examining only the physically reasonable positive lag domain and applying stringent cuts to minimise the effect of aliasing on each individual source. This produces a more restricted sample with fundamentally different statistical properties, which can be seen in the tension between the SDSS and OzDES constraints. To make joint fits to the relation using both the SDSS and OzDES data sets, we required a set of lags with relatively homogeneous selection properties. To that end, we derive an alternate set of lags from the published SDSS data of SDSS-Shen_2023 that differ from their sample in two ways: firstly, in that we use the lags as recovered from JAVELIN instead of their choice of PyROA, and secondly that we apply a number of quality cuts to this alternate sample to bring it more in line with the rationale of the OzDES selections.
@@
For step 2, in cases where the JAVELIN lag peak is strongly constrained, we replace the JAVELIN confidence interval in this test with days for H or days for CIV and MgII, following a similar reasoning to the cuts of OzDES-Penton_2021. After applying these cuts, we retain a sub-sample with H lags, MgII lags and CIV lags from the SDSS sample. JAVELIN’s MCMC sampler is known to exhibit strong aliasing artefacts in seasonal signals (see McDougall_2025_LITMUS), meaning that its use by itself can give misleading results. However, JAVELIN’s use in concert with specifically tuned quality cuts produces a set of reliable lags with a low FPR, as validated by the simulations of OzDES-Penton_2021. We use SDSS’s JAVELIN lags rather than their PyROA lags, as the OzDES quality cuts are tuned specifically for JAVELIN’s systematics.
The motivation of this sub-sampling is to sacrifice much of the SDSS sample’s statistical power by cutting many sources, including discarding what is likely a large fraction of true positives, to yield a smaller sample that has OzDES’s high reliability on a per-source basis. This SDSS-JAVELIN subsample is broadly similar to the published SDSS sample for the H and CIV lags, but prefers a markedly for lower scatter in the MgII relationship. For the MgII lags, these selection criteria drastically reduce the scatter of the lags about their best-fit relationship, and marginally decrease its preferred slope (see Figure 12(a) for a comparison).
@@ 5.1 Lag measurements
In total, the OzDES reverberation mapping program has produced high quality AGN lags, consisting of H lags (OzDES-Malik_2023), MgII lags (OzDES-Yu_2021; OzDES-Yu_2023) and CIV lags (OzDES-Penton_2025). For each of these sources, we calculate the mass using equation 1 and accretion rate using equation 4, where the monochromatic luminosity uses bolometric corrections of Runnoe_2012 (Table 2 in their paper) and we use the virial factor of from SDSS-Shen_2023, as well as their dex inherent scatter. Rest-frame lags range from 333 being days, here and elsewhere in the paper. to , sitting within the range of lags recovered by other works. We find that the distribution of lags over redshift is reasonably consistent with existing reverberation mapping results from the literature (Figure 5, upper left). We note that the distribution of lags changes abruptly as one moves between lines, most notably between the MgII and CIV lines, indicating different lags, and thus radii, for different ions. This is quantified in Section 7. These RM results, as well as the redshifts and luminosities for the relevant sources, are listed in Table LABEL:tab:Results_All_Hbeta for our H sources, Table LABEL:tab:Results_All_MgII for MgII sources, and Table LABEL:tab:Results_All_CIV for CIV sources.
@@ 5.2 Black hole masses from direct RM measurements
We broadly see an increase in mass with redshift (Figure 5 lower left), which is expected, as the AGN at higher redshifts are selected to be more luminous (see, for example, Figure 1 in OzDES-Malik_2024), and this selection effect will dominate any evolutionary effects. We similarly see the expected general increase of more luminous AGN being more massive, but are unable to infer a clear power-law relationship (Figure 5 lower right).
It should be noted also that the estimated accretion rates for high redshift CIV sources can be extremely high (Figure 5 upper right), with most direct RM samples sitting above the threshold used by Du_2015 as the division between low and high accretion sources. In Section 6.1.3 we note an apparent flattening of the CIV curve at high luminosity, and, given that SEAMBH associate high accretion rates with low lags (HBETA_U_2022), this offers a possible physical explanation. Within the OzDES sample, there is a weak trend for more luminous, higher-redshift sources to accrete more strongly (black data points, Figure 5 upper right).
@@ 6 Comparison of New Results With Past Surveys & Tuning of Relationship
Here we examine the relationship using the full array of lags available at the close of this latest generation RM surveys. In addition to lags from OzDES and the final SDSS release of SDSS-Shen_2023, we use a wide array of past H, MgII and CIV lags as outlined in Section 3.3. In Section 6.1, we use a Bayesian approach to constrain the parameters of the relationship for all three reverberating emission lines, and to identify where the results of different surveys are in tension. For each line, we then select a ‘primary’ set of data for constraining an relationship that is most representative of the high redshift results of OzDES and SDSS, with the aim of demonstrating that such relations, marginalised over the footprint of such industrial scale surveys, are now within reach.
@@ 6.1 Comparison of Parameter Constraints
In this section examine what constraints can be places on the relationship for all lines using OzDES data alone, and also using OzDES data in combination with data from other sources. Though OzDES and SDSS have similar physical target populations, the same can not be said for the wealth of literature sources. Further still, different surveys have different lag measurement methods and quality cut criteria, meaning we cannot assume them all to trace the same population in the plane. We resolve some of this issue by using the SDSS data products that most closely match the published OzDES population (see Section 4.2.4), but for other surveys we take a high level approach of only allowing data sets to be used in combination if there is some reasonable combination of parameters that satisfy all surveys in the grouping simultaneously. The rationale here is that every survey will have complex multi-layered selection effects, but we are here concerned only if these impose enough of a difference in the plane that they are significantly in tension with one another. A more complex treatment would require modelling of every survey, their selection effects and pipelines and the AGN population, while the approach taken here is a natural extension of our comparison of the surveys in the plane. In Table 2 we list the parameters for our chosen ‘primary’ data sets, while the following sub-sections elaborate on alternate relationships with different combinations of literature data and quality cuts.
@@
Here, represents the choice of luminosity units, typically or erg/s, and is chosen to minimise the covariance between offset and slope in the posterior distribution. varies with the choice of time units for the lag, and we here follow the convention of all lags being in days.
@@
All fits are performed using the Bayesian analysis tool NumPyro (numpyro), specifically its No U-Turn (NUTS) sampler, a Hamiltonian Monte Carlo (HMC) method that is well suited to unimodal distributions. We note that this fitting is relatively insensitive to the choice of sampler, as it is unimodal and produces approximately Gaussian contours (e.g. Figure 12(a)).
| Line | Luminosity Unit | Slope () | Offset () | Scatter () | Slope - Offset Fit Correlation () |
| H | |||||
| MgII | |||||
| CIV | |||||
@@ 6.1.1 H Relationship
Figure 12(a) (a) shows relationship parameters for data from OzDES, SDSS, the Bentz Collection, the SEAMBH Survey, the LAMP survey, and a number of other miscellaneous sources. Shown also are the contours for SDSS lags as recovered from JAVELIN instead of PyROA, and selected according to the criteria outlined in Section 4.2.4. The constraining power is highest for surveys with a high source-count, like SDSS & SEAMBH, or a wide range of luminosities as in the Bentz Collection.
@@
Figure 12(a) (b) shows the constraints on relationship parameters when combining all datasets that are not in tension with one another. Four such groups are possible:
@@
Where we use a ‘primary’ H grouping elsewhere in the paper, it refers to the first of these groupings. We make this choice of grouping as it combines the data of OzDES and SDSS with a broad set of literature lags, while excluding the lags of the SEAMBH survey that are, by the survey’s design, deliberately intended to be atypical of the broader AGN population.
@@ 6.1.2 MgII Relationship
This yields two mutually consistent groupings, making use of the full SDSS data set and excluding OzDES or including OzDES but restricting the SDSS data. Owing to the complementary correlations of the data sets, both choices have near identical results for the main properties of the relation, but with a slightly higher scatter for the SDSS-based results (Figure 12(a) (c)). We proceed with the OzDES + SDSS-JAVELIN + Zajacek fit as our primary MgII relation as it is derived from the most diverse set of sources while still showing very low scatter about the mean relationship, suggestive of a lower false positive rate.
@@ 6.1.3 Relationship
The tension between the slope of the high and low redshift is manifest in an apparent ‘levelling out’ of the relation at high redshift (Figure 6, bottom panel). This is easiest to see when plotting the residuals against the ‘high-’ fit (our ‘grouping 1)’). As shown in Figure 12, high luminosity / redshift sources give unmistakably and consistently higher lags than the lower luminosity / redshift data.
This nonlinear relationship for CIV may have a physical basis, such as distant AGN having higher accretion rates, which have been associated with lower lags (e.g. Du_2015; HBETA_Hu_2021). A simpler explanation lies in describing this as a selection effect: more luminous AGN have longer rest-frame lags, stretched further at high redshift by time dilation, and are therefore more likely to have lags that exceed the length of the survey. This induces a selection effect in which high redshift / luminosity sources are biased downwards from the true population distribution as the finite survey length imposes a ceiling on what observer-frame lags can be recovered. CIV sources are often the most distant, and consequently the most luminous and associated with the longest observer-frame lags, and so their observations are the most impacted by this effect.
@@
| Survey | Highest Observer Frame Lag (days) | Spectroscopic Baseline (yrs) | Nominal Lag Cutoff (days) |
| OzDES | |||
| SDSS-JAVELIN | |||
@@
This model significantly relaxes the constraints from each survey, introducing the freedom for an un-observed fraction of sources to exist above the lag cutoff. This relaxes the tension between the surveys, creating a single parameter-space region of overlap between all four data sources (Figure 13, top panel). The high luminosity literature sources, drawn from the high- Kaspi sample, are dominated by this observational window effect, loosening their constraining power such that this fit matches the constraints of the more generic fit achieved when excluding the high- literature lags all together (Figure 13, bottom panel). It is this fit, with all CIV data accounted for, that we use as our primary fit for single-epoch mass estimates in Section 8.